OpenStudy (inowalst):

Ellie is dropping marbles into a box one at a time in the following order: red, white, white, blue, blue, blue; red, white, white, blue, blue, blue; ... How many marbles will be in the box right after the 100th blue one is put in?

OpenStudy (gorv):

first consider a small box 1 with red white white blue blue blue just six marble in small box 1 .....we are putting these small boxes in larger one

OpenStudy (gorv):

now in each box there are 3 blue marble

OpenStudy (gorv):

considering 3 in each small box how many small box are needed to make 100 ??

OpenStudy (inowalst):

Divide 100/3 ?

OpenStudy (texaschic101):

look at it this way...for every 6 marbles, 3 of them are blue

OpenStudy (gorv):

yeah that means 33 complete small box + 1 box is in complete

OpenStudy (gorv):

in 33 small boxes blue marbles = 33*3 =99

OpenStudy (gorv):

so first marble in next small box will be 100th blue agree??

OpenStudy (inowalst):

Yes.

OpenStudy (gorv):

in 33 box each have total 6 marble

OpenStudy (gorv):

so in 33 box total number of marble =???

OpenStudy (inowalst):

33*6?

OpenStudy (texaschic101):

yes

OpenStudy (gorv):

yes

OpenStudy (inowalst):

I meant.. 198.

OpenStudy (gorv):

99 blue and total 33*6 marble

OpenStudy (gorv):

now in next box first we will put which marble ???

OpenStudy (gorv):

before blue ??

OpenStudy (inowalst):

white

OpenStudy (gorv):

198+1=199

OpenStudy (gorv):

after white ??

OpenStudy (inowalst):

Oh, blue. Sorry.

OpenStudy (texaschic101):

no...what comes before white

OpenStudy (inowalst):

Red..

OpenStudy (gorv):

red +white+white ....do we put theses before blue ????

OpenStudy (inowalst):

Yes.

OpenStudy (gorv):

so total = 198+1 red+ 2white=198+3=201

OpenStudy (texaschic101):

so there is 3 marbles before the last blue one

OpenStudy (inowalst):

Okay.

OpenStudy (gorv):

so now we have total 201 but 99 blue so we need to put one blue agree??

OpenStudy (gorv):

to make blue = 100

OpenStudy (inowalst):

Yes.

OpenStudy (gorv):

so total marble after 100th blue =??

OpenStudy (inowalst):

301?

OpenStudy (gorv):

in 201 we have 99 blue and we are just adding only one more blue

OpenStudy (texaschic101):

201 marbles + 1 blue marble =

OpenStudy (gorv):

201+1blue =??

OpenStudy (inowalst):

202.

OpenStudy (texaschic101):

BINGO :)

OpenStudy (gorv):

u did it :)

OpenStudy (texaschic101):

very good explaining gorv :)

OpenStudy (inowalst):

Alright then. Thank you so much to both of you!!

OpenStudy (gorv):

thankuu buddy @texaschic101 :)